OpenStudy (anonymous):

A golf club manufacturer makes drivers with 4 different shaft lengths, 3 different lofts, 2 different grips, and 2 different club head materials. How many different combinations are possible?

OpenStudy (ash2326):

No. of drivers or no. of combinations possible= product of no. of ways of choosing one from each = (choosing one shaft length from four)*(choosing one loft from three)*(choosing one grip from two)*(Choosing one type of club material from two) Do you get the idea?

OpenStudy (anonymous):

so is it 0

OpenStudy (ash2326):

Nope, Tell me no. of ways of choosing 1 item from 3?

OpenStudy (anonymous):

I really don t get it

OpenStudy (ash2326):

Do you know what's \(^aC_b\)?

OpenStudy (anonymous):

Nope/I am Really bad at math

OpenStudy (ash2326):

@1dvanessah I gotta bounce now, @Zarkon @saifoo.khan please help here

OpenStudy (ash2326):

Sorry :(

OpenStudy (saifoo.khan):

@satellite73

OpenStudy (anonymous):

yeah not n choose k in this case. counting principle says multiply all choices together

OpenStudy (anonymous):

i.e. \(4\times 3\times 2\times 2\)

OpenStudy (anonymous):

It can't be 48,i multply them(or can it_

OpenStudy (dumbcow):

yep just multiply all the number choices together to get total possible combinations

OpenStudy (anonymous):

looks like 48 to me

OpenStudy (anonymous):

Thanks(: